• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Sirca

    Log entry time 23:11:47 on March21,2005

    Entry number 142139

    This entry is a followup to: 142138

    Followups:

    keyword=I_beam 20 uA to 30 uA

    The vacuum improved significantly, so try to raise I_beam to 30 uA
    and thus achieve stable equilibrium. The other option is to run
    at 40 uA for one hour, let pressure gauges trip the beam, pump down,
    continue 40 uA running, etc.

    This is obivously not a nice way to go on until morning, so MCC
    will try to figure out what is going on while running at 30 uA.
    Beam mis-steering or RF instabilities could be the cause.